# A PAH deficit in the starburst core of a distant spiral galaxy

### v1

@@ Abstract
While PAH features exhibit a tight correlation with galaxy-wide SFR, whether they are reliable tracers of spatially-resolved star formation is still under investigation. For example, PAH molecules can be dissociated by UV photons from star formation or AGN accretion disk as shown by theory (Leger et al., 1989). Using the Spitzer Space Telescope, many studies have reported a deficit of PAH emission in the cores of active galactic nuclei (AGNs) or compact starbursts based on infrared spectrograph (IRS) spectroscopy (e.g. Desai et al., 2007; Díaz-Santos et al., 2011; Stierwalt et al., 2013, 2014; Zhang et al., 2022). Inami et al. (2013) also present similar observations in a survey of 202 local luminous infrared galaxies via the AKARI telescope. Recently, studies using mid-infrared instrument imaging (MIRI) aboard JWST have found evidence of PAH deficits in HII regions in nearby AGNs (e.g., Lai et al., 2022, 2023) and even in normal SFGs (e.g., Egorov et al., 2023; Spilker et al., 2023; Pedrini et al., 2024). Therefore, such deficits could also occur due to strong radiation fields in star-forming regions and their effects should be considered when utilizing PAH emissions to infer SFRs.
@@ 3.2 Spatially-resolved map
Thanks to the high sensitivity and resolution of ALMA and JWST, investigate this relationship on a spatially-resolved manner. The F1800W filter covers the rest-frame wavelength of m of PACS-830, as shown in Figure 3. This allows us to use the flux in F1800W as a spatially resolved proxy for at 0.59′′ resolution. To account for the difference between the flux in F1800W and IRAC- passband, we apply a k-correction using the SED template of M82 from Polletta et al. (2007) to make a fair comparison with the results in Elbaz et al. (2011). The corrected flux is 1.3 times lower. We infer the spatially resolved map from the F1800W image by converting the corrected flux to luminosity. To assess the spatially resolved , we utilize high-resolution ALMA observations of CO (=5–4) instead of the dust continuum. This approach avoids the extra uncertainty introduced by temperature variations in the dust component, as CO (=5–4) flux is independent of dust modeling, allowing us to infer in each pixel at 0.3 arcsec resolution. The conversion is based on the empirical linear relation between CO (=5–4) luminosity and the total infrared luminosity , well demonstrated in Greve et al. (2014); Daddi et al. (2015); Liu et al. (2015); Valentino et al. (2020). We adopt the equation in Daddi et al. (2015), as follows:
@@ 4.1 The multi-wavelength morphology of PACS-830
We emphasize that the CO (=5–4) emission distinctly reveals the spiral arms333Complementary kinematic information from the moment 1 and 2 maps is provided in Appendix B. , with notably clearer detection in the eastern arm. We highlight this by overlaying the CO(5-4) contours on the F444W image, which traces the stellar component (Figure 1.g). We fit the CO (=5–4) emission of the eastern arm and the core simultaneously with two Gaussian models using the Cube Analysis and Rendering Tool for Astronomy (CARTA, Comrie et al., 2021) and find the flux contribution of the arm to the total is 216%.
@@
The SFR map (Figure 2) from pixel-by-pixel SED fitting based on the HST and NIRCam images resembles the CO (=5–4) map and F1800W image, all tracing regions of enhanced star-forming activity, including the galactic core and bright knots to the southeast and north.While the qualitative morphological similarity is evident, we caution against interpreting the SFR map as a quantitative tracer of obscured star formation. It is based solely on rest-frame UV-to-NIR data and assumes a constant star formation history, without incorporating mid- or far-infrared constraints. Thus, variations in the CO/SFR ratio larger than a factor of 2 are expected, and we use the SFR map only as a complementary view of the morphology of star formation regions rather than for quantitative analysis.
The morphology of the starburst core in CO (=5–4) appears distorted, possibly associated with the surrounding components revealed by JWST in Figures 1.l and 2.b. The bulk of star formation is concentrated in the galactic center, indicating a bulge being fueled, which is commonly observed at cosmic noon (e.g., Jiménez-Andrade et al., 2019; Tan et al., 2024a). The dense gas in the southeastern spiral arm appears highly extended in Figure 1.j. The morphological similarity between the pixel-by-pixel SED-fitting SFR map and the CO (=5–4) emission in this region suggests that the extension may consist of two bright clumps blending into a single structure—consistent with how the two clumps in Figure 1.f appear merged in Figure 1.h. To assess the impact of resolution, we downgraded the F277W image to the F770W resolution by PSF matching. The resulting flux ratio map shows no clear evidence of multiple clumps in the arms or bulge, indicating that the "extension" appearance in F770W is largely due to resolution blending. This blending likely contributes to the irregular morphology seen in CO (5–4), where clumps and bulge structures become spatially mixed. Although the precise CO (=5–4) morphology cannot be definitively established given ALMA data’s sensitivity and resolution, the SFR map generated by pixel-by-pixel SED fitting indicates the presence of tentative structure in areas with low signal-to-noise in CO (=5–4).
@@ 4.2 An enhancement in the core: A PAH deficit in the starbursting core?
is defined as the ratio of the infrared luminosity (), to the luminosity at 8 m, , which Elbaz et al. (2011) found to closely correlate with the surface density of MIR and FIR luminosity. In our case, we utilize the CO (=5–4) line as a proxy for and the flux in F1800W as . The uncertainty of is mainly contributed by the conversion from the luminosity of CO (J=5–4) to between , which is a small scatter of 0.2 dex. In Figure 3, the map peaks at the center, with a value around 18, which is roughly double the values observed in the surrounding spiral arms, as indicated by the CO contours. While we acknowledge that our measured may be affected by systematic uncertainties related to CO excitation or temperature variations across different regions of the galaxy, the F1800W data provide a distinct morphological signature: the southeastern spiral arm exhibits a peak flux comparable to the central region. In contrast, the other observations—namely, the SFR map from pixel-by-pixel SED fitting, the CO(=5–4) emission, and the dust continuum—all show a more concentrated central component and a comparatively weaker southeastern spiral arm. This discrepancy suggests that systematic uncertainties alone are unlikely to account for the pronounced variations.
@@
However, some observations find no obvious correlation between the hardness of the radiation field or the emission-line width and IR8 (Inami et al., 2013; Stierwalt et al., 2014). Alternatively, the difference across various star-forming regions may be linked to dust distribution and temperature, as explained by the concept of "dust-bounded" regions (Abel et al., 2009; Díaz-Santos et al., 2017). Dust can absorb a fraction of the ionizing radiation, preventing it from reaching the photodissociation region (PDR). This absorption results in an increase in dust temperature and a decrease in PAH emission from PDRs. In the relation between and , remains relatively flat with respect to the radiation field indicator until it reaches a critical density, , as shown in Figure 4. Beyond this threshold, increases with , a trend confirmed by observations from Elbaz et al. (2018). In PACS-830, we apply two r=0. apertures to measure the in the spiral arm and starbursting core from the original CO (=5–4) map, respectively. As depicted in Figure 4, the result shows that the spiral arm region is at a lower density (), while the core region is of a higher density, at the critical densities (), corresponding to a higher . The core density is not exceptionally high—i.e., it does not significantly exceed the critical density—as expected for a moderate starburst ( MS) with a radius of 0.5′′, as derived from 2D modeling with JWST and ALMA (cf. compact submillimeter sources; Puglisi et al., 2019, 2021; Tan et al., 2024a). However, the density remains sufficient to reveal distinct spatial variations in , consistent with theoretical expectations and prior observations on global properties (Elbaz et al., 2018).
For local galaxies, also indicates the starburstiness , i.e., the distance above the MS. Therefore, we examine in the core region and spiral arm using the SFR (i.e., CO (=5–4)) and stellar mass maps. The SFR is further inferred from the CO (=5–4) converted map based on the calibration by Hao et al. (2011) and Murphy et al. (2011). Since the CO (=5–4) map has a coarser resolution, we first smooth the stellar mass map by doubling the FWHM of its PSF (0.32′′) and then convolve the CO (=5–4) map to match this new Gaussian-like PSF. We then apply the same apertures as in Figure 4 to measure the based on the MS of Speagle et al. (2014). The results well align with the relation in Elbaz et al. (2011), shown in Figure 5, in contrast to the offset in Figure 4. This indicates that, for a given , is lower than expected, suggesting that the star formation mode in this spiral-starburst galaxy may differ from the compact starbursts triggered by major mergers.
@@ 5.3 Grand-design spiral arms
A stellar mass of and a SFR of 7 locates it as a MS galaxy. The mass ratio of it and PACS-830 is about 1/30, which is below the threshold for a minor merger (e.g., 1/10 in Bottrell et al., 2024). Bottrell et al. (2024) highlights that mini mergers can trigger small enhancements in SFRs and asymmetries, aligning with the moderate starburst and non-axisymmetric morphology of PACS-830. Given that PACS-830 shows only a moderate offset from the MS, being a factor of 4 above the relation, the interaction of this mini merger could trigger the SFR enhancement and the formation of the grand-design spiral arms.
@@ Appendix B Moment 1 and 2 maps
While the 40 km s-1 binning improves the overall S/N, it results in relatively coarse velocity resolution. As a consequence, the detailed kinematic modeling is limited by both this coarse sampling and the relatively low intrinsic velocity dispersion, making it difficult to establish a robust velocity structure.

### v2

@@ Abstract
While PAH features exhibit a tight correlation with galaxy-wide SFR, whether they are reliable tracers of spatially-resolved star formation is still under investigation. For example, PAH molecules can be dissociated by UV photons from star formation or AGN accretion disk as shown by theory (Leger et al., 1989). Using the Spitzer Space Telescope, many studies have reported a deficit of PAH emission in the cores of active galactic nuclei (AGNs) or compact starbursts based on infrared spectrograph (IRS) spectroscopy (e.g. Desai et al., 2007; Díaz-Santos et al., 2011; Stierwalt et al., 2013, 2014; Zhang et al., 2022). Inami et al. (2013) also present similar observations in a survey of 202 local luminous infrared galaxies via the AKARI telescope. Recently, studies using mid-infrared instrument imaging (MIRI) aboard JWST have found evidence of PAH deficits in HII regions in nearby AGNs (e.g., Lai et al., 2022, 2023) and even in normal SFGs (e.g., Egorov et al., 2023; Spilker et al., 2023; Pedrini et al., 2024; Battisti et al., 2025). Therefore, such deficits could also occur due to strong radiation fields in star-forming regions and their effects should be considered when utilizing PAH emissions to infer SFRs.
@@ 3.2 Spatially-resolved map
Thanks to the high sensitivity and resolution of ALMA and JWST, we can investigate this relationship on a spatially-resolved manner. The F1800W filter covers the rest-frame wavelength of m of PACS-830, as shown in Figure 3. This allows us to use the flux in F1800W as a spatially resolved proxy for at 0.59′′ resolution. To account for the difference between the flux in F1800W and IRAC- passband, we apply a k-correction using the SED template of M82 from Polletta et al. (2007) to make a fair comparison with the results in Elbaz et al. (2011). The corrected flux is 1.3 times lower. We infer the spatially resolved map from the F1800W image by converting the corrected flux to luminosity. To assess the spatially resolved , we utilize high-resolution ALMA observations of CO (=5–4) instead of the dust continuum. This approach avoids the extra uncertainty introduced by temperature variations in the dust component, as CO (=5–4) flux is independent of dust modeling, allowing us to infer in each pixel at 0.3 arcsec resolution. The conversion is based on the empirical linear relation between CO (=5–4) luminosity and the total infrared luminosity , well demonstrated in Greve et al. (2014); Daddi et al. (2015); Liu et al. (2015); Valentino et al. (2020). We adopt the equation in Daddi et al. (2015), as follows:
@@ 4.1 The multi-wavelength morphology of PACS-830
We emphasize that the CO (=5–4) emission distinctly reveals the spiral arms333Complementary kinematic information and a discussion of the moment 1 and 2 maps, including a modeling attempt, are provided in Appendix B. , with notably clearer detection in the eastern arm. We highlight this by overlaying the CO (=5–4) contours on the F444W image, which traces the stellar component (Figure 1.g). We fit the CO (=5–4) emission of the eastern arm and the core simultaneously with two Gaussian models using the Cube Analysis and Rendering Tool for Astronomy (CARTA, Comrie et al., 2021) and find the flux contribution of the arm to the total is 216%.
The dense gas in the southeastern spiral arm appears highly extended in Figure 1.j, which could result from two bright clumps being blended into a single structure—consistent with how the two clumps in Figure 1.f appear merged in Figure 1.h. To assess the impact of resolution, we degraded the F277W image to the F770W resolution via PSF matching. The resulting flux ratio map shows no clear evidence for multiple clumps in the arms or bulge, indicating that the extended appearance in F770W is largely due to resolution-induced blending.
As shown in Figure 1.j and k, the bulk of star formation, as indicated by CO and continuum emission, is concentrated in the galactic center, suggesting the presence of a bulge in formation, a feature commonly observed at cosmic noon (e.g., Jiménez-Andrade et al., 2019; Tan et al., 2024a). The slightly distorted morphology of the starburst core in CO(=5–4) may similarly arise from limited resolution and sensitivity, as clumps and bulge structures, distinguishable by color in the RGB image (Figure 1.l), could become spatially blended.
@@
The SFR map (Figure 2) from pixel-by-pixel SED fitting based on the HST and NIRCam images shows regions of enhanced star-forming activity, i.e., the galactic core and the spiral arm to the east. These sites of enhanced SF activity are also co-spatial with the peaks in CO (J=5-4).
@@ 4.2 An enhancement in the core: A PAH deficit in the starbursting core?
is defined as the ratio of the infrared luminosity (), to the luminosity at 8 m, , which Elbaz et al. (2011) found to closely correlate with the surface density of MIR and FIR luminosity. In our case, we utilize the CO (=5–4) line as a proxy for and the flux in F1800W as . The uncertainty of is mainly contributed by the conversion from the luminosity of CO (J=5–4) to between , which is a small scatter of 0.2 dex. In Figure 3, the map peaks at the center, with a value around 18, which is roughly double the values observed in the surrounding spiral arms, as indicated by the CO contours. While we acknowledge that our measured may be affected by systematic uncertainties related to CO excitation or temperature variations across different regions of the galaxy, the F1800W data provide a distinct morphological signature: the southeastern spiral arm exhibits a peak flux approximately 80% that of the central region. In contrast, the CO(=5–4) emission and the dust continuum show more centrally concentrated morphologies with weaker emission in the southeastern arm. This discrepancy suggests that systematic uncertainties alone are unlikely to account for the pronounced variations.
@@
However, some observations find no obvious correlation between the hardness of the radiation field or the emission-line width and IR8 (Inami et al., 2013; Stierwalt et al., 2014). Alternatively, the difference across various star-forming regions may be linked to dust distribution and temperature, as explained by the concept of "dust-bounded" regions (Abel et al., 2009; Díaz-Santos et al., 2017). Dust can absorb a fraction of the ionizing radiation, preventing it from reaching the photodissociation region (PDR). This absorption results in an increase in dust temperature and a decrease in PAH emission from PDRs. In the relation between and , remains relatively flat with respect to the radiation field indicator until it reaches a critical density, , as shown in Figure 4. Beyond this threshold, increases with , a trend confirmed by observations from Elbaz et al. (2018). In PACS-830, we apply three r=0. apertures to measure the in the eastern and northern spiral arm and starbursting core from the original CO (=5–4) map, respectively. We note that, since the northern arm is not robustly detected in CO (=5–4), we treat the CO luminosity as an upper limit. Consequently, the derived values of IR8 and for this region should also be considered upper limits. As depicted in Figure 4, the result shows that the spiral arm regions are at a lower density (), while the core region is of a higher density, at the critical densities (), corresponding to a higher . The core density is not exceptionally high—i.e., it does not significantly exceed the critical density—as expected for a moderate starburst ( MS) with a radius of 0.5′′, as derived from 2D modeling with JWST and ALMA (cf. compact submillimeter sources; Puglisi et al., 2019, 2021; Tan et al., 2024a). However, the density remains sufficient to reveal distinct spatial variations in , consistent with theoretical expectations and prior observations on global properties (Elbaz et al., 2018).
For local galaxies, also indicates the starburstiness , i.e., the distance above the MS. Therefore, we examine in the core region and spiral arms using the SFR (i.e., CO (=5–4)) and stellar mass maps. The SFR is further inferred from the CO (=5–4) converted map based on the calibration by Hao et al. (2011) and Murphy et al. (2011). Since the CO (=5–4) map has a coarser resolution, we first smooth the stellar mass map by doubling the FWHM of its PSF (0.32′′) and then convolve the CO (=5–4) map to match this new Gaussian-like PSF. We then apply the same apertures as in Figure 4 to measure the based on the MS of Speagle et al. (2014). The results well align with the relation in Elbaz et al. (2011), shown in Figure 5, in contrast to the offset in Figure 4. This indicates that, for a given , is lower than expected, suggesting that the star formation mode in this spiral-starburst galaxy may differ from the compact starbursts triggered by major mergers.
@@ 5.3 Grand-design spiral arms
A stellar mass of and a SFR of 7 locates it as a MS galaxy. The mass ratio of it and PACS-830 is about 1/30, which is below the threshold for a minor merger (e.g., 1/10 in Bottrell et al., 2024). Bottrell et al. (2024) highlights that mini mergers can trigger small enhancements in SFRs and asymmetries, aligning with the moderate starburst and non-axisymmetric morphology of PACS-830. Given that PACS-830 shows only a moderate offset from the MS, being a factor of 4 above the relation, the interaction of this mini merger could trigger the SFR enhancement and the formation of the grand-design spiral arms. However, this galaxy is undetected in all ALMA maps and shows only a marginal detection with S/N 4 in F1800W, preventing any further analysis of its PAH and CO emissions.
@@ Appendix B Moment 1 and 2 maps
The moment 1 map reveals a disk-like velocity field, consistent with rotation, and even shows hints of a spiral feature. The moment 2 map also aligns with expectations for a rotating disk, displaying elevated velocity dispersion in the central region. However, we note that the current spatial resolution and signal-to-noise ratio limit a more detailed interpretation. In particular, while the central enhancement in dispersion is qualitatively consistent with the presence of a bulge, it is also consistent with expectations for a disk without requiring a bulge component, unless a proper model comparison is made.
We also attempted to model the CO kinematics using the 3DBarolo code (Teodoro & Fraternali, 2015). While the residual maps were generally acceptable, the fitted dispersion values carried large uncertainties and exceeded the measured values. This may reflect limitations in the modeling or a relatively low intrinsic velocity dispersion (10 km/s). Furthermore, our spectral binning (40 km/s) is significantly larger than the expected dispersion, making the results unreliable. For these reasons, we decided not to include quantitative 3DBarolo results in either the main text or the appendix.